# The Tropical Geometry of Subtraction Schemes

### v1

@@ 8 Conclusion
In this paper we have applied tropical geometry to the study of local subtraction schemes for Euler integrals. Our first main result is a criterion for a combination of Euler integrands to be locally finite, expressed by Th. 3 and Th. 4. We have applied this theorem, and the tropical wisdom that goes into its proof, to construct a simple-minded subtraction scheme for Euler integrals which satisfy the geometrical property 1. The scheme can be summarized by the subtraction formula described in Th. 6, which is the central result of the paper.
@@
Beyond Feynman Integrals The subtraction scheme described in this paper is applicable to any Euler integrand satisfying the geometric property, as it was shown in the provided examples. A particularly satisfying aspect of this framework is that it illuminates the combinatorics of the cancellations. Traditional challenges of subtraction schemes, such as treating double-counting of divergences, are neatly dealt by the face lattice of the Newton polytope. Based on this, we advocate the search for parametrization of integrals appearing in various physical applications, such as cosmology and phase-space integrations, in order to express these quantities in terms of Euler integrand and apply the ideas put forward in this paper.

### v2

@@ 8 Conclusion
In this paper we have applied tropical geometry to the study of local subtraction schemes for Euler integrals. Our first main result is a criterion for a combination of Euler integrands to be locally finite, expressed by Th. 3 and Th. 4. We have applied this theorem, and the tropical wisdom that goes into its proof, to construct a simple-minded subtraction scheme for Euler integrals which satisfy the geometrical property 1. The scheme can be summarized by the subtraction formula described in Th. 6, which is the central result of the paper. The Nilsson-Passare analytical continuation allows to algorithmically express any Euler integral as a combination of integrals satisfying the assumption of Th. 6, so our results apply to arbitrary Euler integrals. A private implementation of this method is available upon request to the author, and it will be made public in the future.
@@
Beyond Feynman Integrals The subtraction scheme described in this paper is applicable to any Euler integrand satisfying the geometric property, as it was shown in the provided examples. A particularly satisfying aspect of this framework is that it illuminates the combinatorics of the cancellations. Traditional challenges of subtraction schemes, such as treating double-counting of divergences, are neatly dealt by the face lattice of the Newton polytope. Based on this, we advocate for the search of parametrizations of integrals appearing in various physical applications, such as cosmology and phase-space integrations, in order to express these quantities in terms of Euler integrals.